the base of a rectangle is 4.8 units shorter than the length h, write an equation,in terms of h, when the area is 150 square units
Answer by Cromlix(4381)

Hi there,
Length = h
Width = h - 4.8
Area = Length x Width
150 = h(h - 4.8)
150 = h^2 - 4.8h
h^2 - 4.8h - 150 = 0
Hope this helps :-)